We cannot use Mathematical Induction to make recursive definitions.

True or False?
Answer #1

False.

For example , suppose the the recursive relation an is defined by ,

an = an-1 + n and a1 = 1

We can use induction on n to prove that ,

an = \frac{1}{2}n (n +1) .

So false.